OpenStudy (anonymous):

convert (1777) base 8 to hexadecimal

OpenStudy (anonymous):

first convert it to decimal no. system 8^0 *7 + 8^1 *7 + 8^2 *7 +8^3 *1 = 7 +56 + 448 + 512 = 1023 now convert it to hexadecimal no. system divide by 6 and remaining remainders will give you the answer your ans is 3244